# MambaMoE: Mixture-of-Spectral-Spatial-Experts State Space Model for Hyperspectral Image Classification

### v1

@@ Abstract.
The Mamba model has recently demonstrated strong potential in hyperspectral image (HSI) classification, owing to its ability to perform context modeling with linear computational complexity. However, existing Mamba-based methods usually neglect the spectral and spatial directional characteristics related to heterogeneous objects in hyperspectral scenes, leading to limited classification performance. To address these issues, we propose MambaMoE, a novel spectral-spatial mixture-of-experts framework, representing the first MoE-based approach in the HSI classification community. Specifically, we design a Mixture of Mamba Expert Block (MoMEB) that leverages sparse expert activation to enable adaptive spectral-spatial modeling. Furthermore, we introduce an uncertainty-guided corrective learning (UGCL) strategy to encourage the model’s attention toward complex regions prone to prediction ambiguity. Extensive experiments on multiple public HSI benchmarks demonstrate that MambaMoE achieves state-of-the-art performance in both accuracy and efficiency compared to existing advanced approaches, especially for Mamba-based methods. Code will be released.
Hyperspectral Image (HSI) Classification, Mixture-of-Experts, Mamba, Spectral-Spatial , Uncertainty
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Hyperspectral imaging
1. Introduction
@@
Hyperspectral images (HSIs) contain hundreds of contiguous spectral bands, enabling the capture of rich spectral-spatial information that supports accurate analysis for Earth observation tasks (Luo et al., 2024; Li et al., 2024a). HSI classification, which involves assigning a semantic label to each pixel in the image, has emerged as a fundamental task in the remote sensing community.
@@
With the rise of deep learning, HSI classification has transitioned from handcrafted feature engineering (Guo et al., 2008; Bengio et al., 2013; Bernard et al., 2012) to fully automated neural network-based approaches. Early methods, such as CNNs (Hu et al., 2015; Zhao and Du, 2016), demonstrated strong local feature extraction capabilities but were limited in capturing global context due to their constrained receptive fields. To address this, Transformer-based models (Hong et al., 2022; Wang et al., 2023) introduced long-range dependency modeling. However, their reliance on the self-attention mechanism (Vaswani et al., 2017; Xu et al., 2021) incurs quadratic computational complexity, making them less efficient for modeling spatial and spectral contexts simultaneously.
@@
Recently, Mamba-based architectures have garnered increasing attention in HSI classification due to their context modeling capability while maintaining linear complexity (Gu et al., 2021; Gu and Dao, 2023). However, existing approaches exhibit several limitations. Firstly, most existing methods adopt patch-level designs (Wang et al., 2025; Pan et al., 2025), where an image patch is fed into the network to predict the category of the central pixel. This approach restricts the receptive field and fails to capture the holistic spatial context of remote sensing scenes. In addition, although multi-directional spectral-spatial scanning (Liu et al., 2024b) is commonly employed to enhance feature diversity, existing methods typically treat all directions equally, overlooking the directional variations inherent in heterogeneous land-cover types (see Fig. 1). Furthermore, the inherently sequential modeling (Liu et al., 2024a; Zhu et al., 2024) paradigm of Mamba disrupts the original spatial structure of hyperspectral data, resulting in a loss of spatial coherence. This degradation limits the model’s ability to preserve and exploit critical spatial information, ultimately hindering the extraction of fine-granularity and context-aware features that are essential for accurate classification.
Figure 1. Illustration of various directional characteristics of heterogeneous objects in HSIs.
@@
To address these limitations, inspired by the Mixture-of-Experts (MoE) paradigm (Shazeer et al., 2017; Riquelme et al., 2021; Puigcerver et al., 2024), we propose MambaMoE—a spectral-spatial mixture-of-mamba-expert framework for HSI classification. This framework is composed of specialized network blocks based on Mamba structures, which leverage state space models oriented along distinct spatial and spectral directions. Within these blocks, multiple Mamba modules serve as diverse experts, collectively forming the proposed Mixture of Mamba Expert Block (MoMEB). Given that remote sensing images are captured from a bird’s-eye view—where objects can appear in arbitrary orientations—and hyperspectral data channels follow a fixed wavelength order, i.e., exhibiting directional spectral dependencies, MoMEB is designed to reflect these spatial-spectral characteristics. Specifically, it treats each spatial-directional Mamba branch as a routed expert and employs two spectral-directional branches as shared experts. Furthermore, a lightweight router network is integrated within MoMEB to dynamically activate a subset of experts for each input, enabling the model to adaptively extract informative spectral-spatial directional features. This design significantly enhances classification performance while maintaining computational efficiency.
@@
Different from existing patch-level Mamba classification networks (Huang et al., 2024; Liu et al., 2024b), MambaMoE is an end-to-end image-level segmentation network built upon a classical encoder-decoder architecture. As described above, MoMEBs are employed in the encoder to extract representative spectral-spatial features. However, due to the limited directional diversity in Mamba’s sequential modeling, these features may not fully capture the varied object orientations present in complex HSI scenes, potentially leading to classification ambiguity. To mitigate this, we introduce uncertainty-guided corrective learning (UGCL) in the decoder. Specifically, UGCL encourages the model to focus on challenging regions by dynamically sampling supervisory labels from areas of high prediction uncertainty, guiding MoMEB outputs to adaptively refine their representations. This strategy enables the network to better handle difficult cases and enhances its generalization capability. Thanks to this design, MambaMoE achieves state-of-the-art performance across multiple public HSI benchmarks, surpassing existing advanced methods, including those based on Mamba architectures.
@@
2. Related work
2.1. Mamba-based HSI Classification
@@
Recently, Mamba has attracted increasing attention in the field of HSI classification due to its ability to achieve context modeling with linear computational complexity (Liu et al., 2024a). Yao et al. (Yao et al., 2024) proposed SpectralMamba, which simplifies sequence learning in the state space and enhances spectral correction within the spectral-spatial domain. SS-Mamba (Huang et al., 2024) further extends this idea by introducing multiple spectral-spatial Mamba modules. In addition, Li et al. (Li et al., 2024b) proposed MambaHSI, which models long-range dependencies across the entire image and adaptively fuses spatial and spectral information. HyperMamba (Liu et al., 2024b) introduces a spectral-spatial adaptive architecture that jointly performs spatial neighborhood scanning and spectral feature enhancement. He et al. (He et al., 2024) explored the application of Mamba from a 3D perspective, focusing on 3D cubes. MambaLG (Pan et al., 2025) employs a dual-branch architecture that integrates local-global spatial modeling and dynamic short- and long-range spectral feature perception, fully leveraging Mamba’s sequence modeling capabilities. Wang et al. (Wang et al., 2025) proposed S2Mamba, a spectral-spatial scanning-based model that employs a learnable mixture gate to effectively fuse spatial and spectral representations. Building upon this line of work, our method also adopts image-level input, as used in MambaHSI (Li et al., 2024b), but further enhances modeling capacity by integrating a sparsely activated MoE in the encoder, enabling dynamic and adaptive spectral-spatial representation. Additionally, we design the UGCL strategy to dynamically guide the model to learn from labels sampled in complex regions, effectively mitigating classification ambiguities in challenging areas.
2.2. Mixture-of-Experts
@@
Mixture-of-Experts (MoE) have emerged as a promising paradigm for enhancing model efficiency and scalability by enabling dynamic, data-dependent routing through specialized sub-networks (experts). MoE has been widely adopted in the NLP community to scale large language models, as exemplified by MoA (Zhang et al., 2022), DeepseekMoE (Dai et al., 2024) and others (Zhou et al., 2022; Fedus et al., 2022). Beyond NLP, MoE has also gained traction in the computer vision domain. For instance, Mod-Squad (Chen et al., 2023) utilizes MoE for efficient multi-task learning, while RingMoE (Bi et al., 2025) adopts a sparse MoE design to build a unified multi-modal foundation model for remote sensing. Despite its success in both NLP and vision tasks, its application to HSI classification remains largely unexplored. In this work, we introduce the MoE into HSI classification for the first time, enabling adaptive spectral-spatial joint feature learning through the proposed spatially routed experts and spectrally shared experts.
@@
Figure 2. An illustration of the proposed MambaMoE. The core component, MoMEB incorporates DSSEM, which leverages SRE and SSE to dynamically capture spectral-spatial features. The Decoder is composed of FFBs and UARBs. After feature extraction and refinement, we adopt the fully connected layer as the classification head to generate the final predictions.
3. Method
3.1. Overview
@@
(1) $\begin{split}&\mathbf{F}_{1}=Avg(ReLU(Conv(\mathbf{X})))\\ &\mathbf{F}_{i}=Avg(ReLU(Conv(\mathbf{F}_{i-1}))),i=2,3.\end{split}$
Here, $Conv(\cdot)$, $GN(\cdot)$, $ReLU(\cdot)$, and $Avg(\cdot)$ denote a $3\times 3$ convolution operation, a ReLU activation function, and an average pooling layer, respectively. Subsequently, these features are fed into the proposed MoMEB modules to adaptively capture spectral-spatial contexts, defined as: $\mathbf{M}_{i}=MoMEB(\mathbf{F}_{i}),\mathbf{M}_{i}\in\mathbb{R}^{C\times H_{i}\times W_{i}},i=1,2,3$. In the decoder, a group of Feature Fusion Blocks (FFBs) are applied to integrate inter-scale features from different stages, which can be formulated as:
(2) $\mathbf{L}_{i}=\begin{cases}\text{FFB}(\mathbf{M}_{i},\mathbf{L}_{i+1}),&i=1,2\\ \text{FFB}(\mathbf{M}_{i}),&i=3\end{cases}$
@@
(3) $\text{Res}(x)=x+\text{Conv2}_{3\times 3}(\text{ReLU}(\text{Conv1}_{3\times 3}(\text{ReLU}(x))))$
@@
3.2. Mixture of Mamba Expert Block
@@
Given an input feature map $\mathbf{F}$ (the stage index is omitted in later texts for convenience), it is first normalized by a LayerNorm (LN) layer, obtaining a new feature $\overline{\mathbf{F}}$. Next, this feature is processed by the Dynamic Spectral-Spatial Expert Mamba (DSSEM) module. Specifically, the input feature is split along the channel dimension into two equal parts, forming a spatial view $\mathbf{F}_{spa}\in\mathbb{R}^{(C/2)\times h\times w}$ and a spectral view $\mathbf{F}_{spe}\in\mathbb{R}^{(C/2)\times h\times w}$, where $h$ and $w$ are height and width of the feature in corresponding stages. These two views are separately fed into the SRE and SSE modules to extract spatial and spectral contextual features, respectively. The resulting outputs, $SRE(\mathbf{F}_{spa})$ and $SSE(\mathbf{F}_{spe})$, are concatenated and then passed through a $1\times 1$ convolution to integrate the information, producing a dynamically modulated spectral-spatial fusion feature $\hat{\mathbf{F}}$. The overall procedure of DSSEM is
(4) $\begin{split}&\mathbf{F}_{spa},\mathbf{F}_{spe}=ChannelSplit(\overline{\mathbf{F}})\\ &\hat{\mathbf{F}}=Conv_{1\times 1}(Concat(SRE(\mathbf{F}_{spa}),SSE(\mathbf{F}_{spe})))\end{split}$
@@
(5) $F_{spa}^{j}=\left\{[f_{j}^{1},f_{j}^{2},\cdots,f_{j}^{h\times w}]\;\middle|\;f_{j}\in\mathbb{R}^{1\times(C/2)},\;j\in\{1,2,3,4\}\right\}$
@@
(6) $\mathbf{h}_{j}^{t}=\overline{\mathbf{A}}^{spa}\mathbf{h}_{j}^{t-1}+\overline{\mathbf{B}}^{spa}f_{j}^{t},$
@@
(7) $\mathbf{w}=\text{softmax}(\mathcal{G}(\mathbf{F}_{spa}))$
@@
(8) $\mathbf{\hat{F}}_{spa}=\sum_{j=1}^{4}\mathbf{w}_{j}\cdot\mathcal{E}_{j}(\mathbf{F}_{spa}).$
@@
Shared Spectral Expert. The SSE module performs essential spectral context modeling and functions as a shared spectral feature extractor, enabling the model to capture consistent spectral patterns across different spatial locations. In the SSE module, forward and backward spectral sequences are constructed along the spectral dimension and processed using SSMs to extract bi-directional spectral features. As illustrated in Fig. 2, the input feature $\mathbf{F}_{spe}$ is first flattened along the spatial dimensions to form spectral sequences:
(9) $F^{l}_{spe}=\left\{[\tilde{f}_{l}^{0},\tilde{f}_{l}^{1},\cdots,\tilde{f}_{l}^{(C/2)}]\;\middle|\;\tilde{f}_{l}\in\mathbb{R}^{1\times(h\cdot w)},\;l\in\{1,2\}\right\}.$
@@
(10) $\mathbf{\tilde{h}}_{l}^{t}=\overline{\mathbf{A}}^{spe}\mathbf{\tilde{h}}_{l}^{t-1}+\overline{\mathbf{B}}^{spe}\tilde{f}_{l}^{t},$
@@
Figure 3. The overview of the UARB.
3.3. Uncertainty-Guided Corrective Learning
In addition to the MoMEBs, we further enhance predictive performance in challenging regions by encouraging the model to learn from dynamically identified difficult areas, as determined by uncertainty estimation. To this end, we propose the UARB, as illustrated in Fig. 3. The UARB processes intermediate features $\mathbf{L}\in\mathbb{R}^{C\times h\times w}$ from various decoder stages.

### v2

@@ Abstract
Mamba-based models have recently demonstrated significant potential in hyperspectral image (HSI) classification, primarily due to their ability to perform contextual modeling with linear computational complexity. However, existing Mamba-based approaches often overlook the directional modeling heterogeneity across different land-cover types, leading to limited classification performance. To address these limitations, we propose MambaMoE, a novel spectral-spatial Mixture-of-Experts (MoE) framework, which represents the first MoE-based approach in the HSI classification domain. Specifically, we design a Mixture of Mamba Expert Block (MoMEB) that performs adaptive spectral-spatial feature modeling via a sparse expert activation mechanism. Additionally, we introduce an uncertainty-guided corrective learning (UGCL) strategy that encourages the model to focus on complex regions prone to prediction ambiguity. This strategy dynamically samples supervision signals from regions with high predictive uncertainty, guiding the model to adaptively refine feature representations and thereby enhancing its focus on challenging areas. Extensive experiments conducted on multiple public HSI benchmark datasets show that MambaMoE achieves state-of-the-art performance in both classification accuracy and computational efficiency compared to existing advanced methods, particularly Mamba-based ones. The code will be available online at https://github.com/YichuXu/MambaMoE.
keywords:
[…]
1 Introduction
@@
Hyperspectral images (HSIs) contain hundreds of contiguous spectral bands, enabling the capture of rich spectral-spatial information that supports accurate analysis for Earth observation tasks [1, 2, 3, 4, 5]. HSI classification, which involves assigning a semantic label to each pixel in the image, has emerged as a fundamental task in the remote sensing community [6, 7, 8].
@@
With the rise of deep learning, HSI classification has transitioned from handcrafted feature engineering methods, such as tensor singular spectrum analysis [9, 10, 11], to fully automated neural network-based approaches. Early methods, such as CNNs [12, 13], demonstrated strong local feature extraction capabilities but were limited in capturing global context due to their constrained receptive fields. To address this, Transformer-based models [14, 15, 16] introduced long-range dependency modeling. However, their reliance on the self-attention mechanism [17, 18] incurs quadratic computational complexity, making them less efficient for modeling spatial and spectral contexts simultaneously.
@@
Recently, Mamba-based architectures have garnered increasing attention in HSI classification due to their context modeling capability while maintaining linear complexity [19, 20]. However, existing approaches exhibit several limitations. Firstly, most existing methods adopt patch-level designs [21, 22], where an image patch is fed into the network to predict the category of the central pixel. This approach restricts the receptive field and fails to capture the holistic spatial context of remote sensing scenes. In addition, although multi-directional spectral-spatial scanning [23] is commonly employed to enhance feature diversity, existing methods typically treat all directions equally, overlooking the directional variations inherent in heterogeneous land-cover types (see Section 4.5). Furthermore, the inherently sequential modeling [24] paradigm of Mamba disrupts the original spatial structure of hyperspectral data, resulting in a loss of spatial coherence. This degradation limits the model’s ability to preserve and exploit critical spatial information, ultimately hindering the extraction of fine-granularity and context-aware features that are essential for accurate classification.
@@
To address these limitations, inspired by the Mixture-of-Experts (MoE) paradigm [25, 26], we propose MambaMoE—a spectral-spatial mixture-of-mamba-expert framework for HSI classification. This framework is composed of specialized network blocks based on Mamba structures, which leverage state space models oriented along distinct spatial and spectral directions. Within these blocks, multiple Mamba modules serve as diverse experts, collectively forming the proposed Mixture of Mamba Expert Block (MoMEB). Given that remote sensing images are captured from a bird’s-eye view—where objects can appear in arbitrary orientations—and hyperspectral data channels follow a fixed wavelength order, i.e., exhibiting directional spectral dependencies, MoMEB is designed to reflect these spatial-spectral characteristics. Specifically, it treats each spatial-directional Mamba branch as a routed expert and employs two spectral-directional branches as shared experts. Furthermore, a lightweight router network is integrated within MoMEB to dynamically activate a subset of experts for each input, enabling the model to adaptively extract informative spectral-spatial directional features. This design significantly enhances classification performance while maintaining computational efficiency.
@@
Different from existing patch-level Mamba classification networks [27, 23], MambaMoE is an end-to-end image-level segmentation network built upon a classical encoder-decoder architecture. As described above, MoMEBs are employed in the encoder to extract representative spectral-spatial features. However, due to the limited directional diversity in Mamba’s sequential modeling, these features may not fully capture the varied object orientations present in complex HSI scenes, potentially leading to classification ambiguity. To mitigate this, we introduce uncertainty-guided corrective learning (UGCL) in the decoder. Specifically, UGCL encourages the model to focus on challenging regions by dynamically sampling supervisory labels from areas of high prediction uncertainty, guiding MoMEB outputs to adaptively refine their representations. This strategy enables the network to better handle difficult cases and enhances its generalization capability. Thanks to this design, MambaMoE achieves state-of-the-art performance across multiple public HSI benchmarks, surpassing existing advanced methods, including those based on Mamba architectures.
@@
The remaining section of this paper is organized as follows. Section 2 reviews the related work. Section 3 provides a comprehensive description of the proposed MambaMoE model. Section 4 provides a comprehensive presentation and analysis of the experimental results. Finally, Section 5 concludes the paper.
2 Related Work
2.1 Mamba-based HSI Classification
@@
Recently, Mamba has attracted increasing attention in the field of HSI classification due to its ability to achieve context modeling with linear computational complexity. Yao et al. [28] proposed SpectralMamba, which simplifies sequence learning in the state space and enhances spectral correction within the spectral-spatial domain. SS-Mamba [27] further extends this idea by introducing multiple spectral-spatial Mamba modules. In addition, Li et al. [29] proposed MambaHSI, which models long-range dependencies across the entire image and adaptively fuses spatial and spectral information. HyperMamba [23] introduces a spectral-spatial adaptive architecture that jointly performs spatial neighborhood scanning and spectral feature enhancement. He et al. [30] explored the application of Mamba from a 3D perspective, focusing on 3D cubes. MambaLG [22] employs a dual-branch architecture that integrates local-global spatial modeling and dynamic short- and long-range spectral feature perception, fully leveraging Mamba’s sequence modeling capabilities. Wang et al. [21] proposed S2Mamba, a spectral-spatial scanning-based model that employs a learnable mixture gate to effectively fuse spatial and spectral representations. Building upon this line of work, our method also adopts image-level input, as used in MambaHSI [29], but further enhances modeling capacity by integrating a sparsely activated MoE in the encoder, enabling dynamic and adaptive spectral-spatial representation. Additionally, we design the UGCL strategy to dynamically guide the model to learn from labels sampled in complex regions, effectively mitigating classification ambiguities in challenging areas.
2.2 Mixture-of-Experts
@@
Mixture-of-Experts (MoE) have emerged as a promising paradigm for enhancing model efficiency and scalability by enabling dynamic, data-dependent routing through specialized sub-networks (experts). MoE has been widely adopted in the NLP community to scale large language models, as exemplified by MoA [31], DeepseekMoE [32] and others [33]. Beyond NLP, MoE has also gained traction in the computer vision domain. For instance, Mod-Squad [34] utilizes MoE for efficient multi-task learning, while RingMoE [26] adopts a sparse MoE design to build a unified multi-modal foundation model for remote sensing. Despite its success in both NLP and vision tasks, its application to HSI classification remains largely unexplored. In this work, we introduce the MoE into HSI classification for the first time, enabling adaptive spectral-spatial joint feature learning through the proposed spatially routed experts and spectrally shared experts.
3 Methodology
3.1 Overview
@@
Figure 1: An illustration of the proposed MambaMoE. The core component, MoMEB incorporates DSSEM, which leverages SRE and SSE to dynamically capture spectral-spatial features. The Decoder is composed of FFBs and UARBs. After feature extraction and refinement, we adopt the fully connected layer as the classification head to generate the final predictions.
@@
$\begin{split}&\mathbf{F}_{1}=Avg(ReLU(Conv(\mathbf{X})))\\ &\mathbf{F}_{i}=Avg(ReLU(Conv(\mathbf{F}_{i-1}))),i=2,3.\end{split}$ (1)
Here, $Conv(\cdot)$, $GN(\cdot)$, $ReLU(\cdot)$, and $Avg(\cdot)$ denote a $3\times 3$ convolution operation, a ReLU activation function, and an average pooling layer, respectively. Subsequently, these features are fed into the proposed MoMEB modules to adaptively capture spectral-spatial contexts, defined as: $\mathbf{M}_{i}=\text{MoMEB}(\mathbf{F}_{i}),\mathbf{M}_{i}\in\mathbb{R}^{C\times H_{i}\times W_{i}},i=1,2,3$. In the decoder, a group of Feature Fusion Blocks (FFBs) are applied to integrate inter-scale features from different stages, which can be formulated as:
$\mathbf{L}_{i}=\begin{cases}\text{FFB}(\mathbf{M}_{i},\mathbf{L}_{i+1}),&i=1,2\\ \text{FFB}(\mathbf{M}_{i}),&i=3\end{cases}$ (2)
@@
$\text{Res}(x)=x+\text{Conv2}_{3\times 3}(\text{ReLU}(\text{Conv1}_{3\times 3}(\text{ReLU}(x))))$ (3)
@@
3.2 Mixture of Mamba Expert Block
@@
Given an input feature map $\mathbf{F}$ (the stage index is omitted in later texts for convenience), it is first normalized by a LayerNorm (LN) layer, obtaining a new feature $\overline{\mathbf{F}}$. Next, this feature is processed by the Dynamic Spectral-Spatial Expert Mamba (DSSEM) module. Specifically, the input feature is split along the channel dimension into two equal parts, forming a spatial view $\mathbf{F}_{spa}\in\mathbb{R}^{(C/2)\times h\times w}$ and a spectral view $\mathbf{F}_{spe}\in\mathbb{R}^{(C/2)\times h\times w}$, where $h$ and $w$ are height and width of the feature in corresponding stages. These two views are separately fed into the SRE and SSE modules to extract spatial and spectral contextual features, respectively. The resulting outputs, $\text{SRE}(\mathbf{F}_{spa})$ and $\text{SSE}(\mathbf{F}_{spe})$, are concatenated and then passed through a $1\times 1$ convolution to integrate the information, producing a dynamically modulated spectral-spatial fusion feature $\hat{\mathbf{F}}$. The overall procedure of DSSEM is
$\begin{split}&\mathbf{F}_{spa},\mathbf{F}_{spe}=\text{ChannelSplit}(\overline{\mathbf{F}})\\ &\hat{\mathbf{F}}=Conv_{1\times 1}(\text{Concat}(\text{SRE}(\mathbf{F}_{spa}),\text{SSE}(\mathbf{F}_{spe})))\end{split}$ (4)
@@
$F_{spa}^{j}=\left\{[f_{j}^{1},f_{j}^{2},\cdots,f_{j}^{h\times w}]\;\middle|\;f_{j}\in\mathbb{R}^{1\times(C/2)},\;j\in\{1,2,3,4\}\right\}$ (5)
@@
$\mathbf{h}_{j}^{t}=\overline{\mathbf{A}}^{spa}\mathbf{h}_{j}^{t-1}+\overline{\mathbf{B}}^{spa}f_{j}^{t},$ (6)
@@
$\mathbf{w}=\text{softmax}(\mathcal{G}(\mathbf{F}_{spa}))$ (7)
@@
$\mathbf{\hat{F}}_{spa}=\sum_{j=1}^{4}\mathbf{w}_{j}\cdot\mathcal{E}_{j}(\mathbf{F}_{spa}).$ (8)
@@
Spectral Shared Expert. The SSE module performs essential spectral context modeling and functions as a shared spectral feature extractor, enabling the model to capture consistent spectral patterns across different spatial locations. In the SSE module, forward and backward spectral sequences are constructed along the spectral dimension and processed using SSMs to extract bi-directional spectral features. As illustrated in Fig. 1, the input feature $\mathbf{F}_{spe}$ is first flattened along the spatial dimensions to form spectral sequences:
$F^{l}_{spe}=\left\{[\tilde{f}_{l}^{0},\tilde{f}_{l}^{1},\cdots,\tilde{f}_{l}^{(C/2)}]\;\middle|\;\tilde{f}_{l}\in\mathbb{R}^{1\times(h\cdot w)},\;l\in\{1,2\}\right\}.$ (9)
@@
$\mathbf{\tilde{h}}_{l}^{t}=\overline{\mathbf{A}}^{spe}\mathbf{\tilde{h}}_{l}^{t-1}+\overline{\mathbf{B}}^{spe}\tilde{f}_{l}^{t},$ (10)
@@
Figure 2: The overview of the UARB.
3.3 Uncertainty-Guided Corrective Learning
In addition to the MoMEBs and inspired by [35], we further enhance predictive performance in challenging regions by encouraging the model to learn from dynamically identified difficult areas, as determined by uncertainty estimation. To this end, we propose the UARB, as illustrated in Fig. 2. The UARB processes intermediate features $\mathbf{L}\in\mathbb{R}^{C\times h\times w}$ from various decoder stages.